OpenStudy (anonymous):

Find the rectangular coordinates of the point with the polar coordinates. -7, 2pi/3

OpenStudy (anonymous):

The relationship between rectangular (x,y) and polar (r,theta) coordinates is x= r*cos(theta) y= r*sin(theta) Ergo a polar coordinate of (-7,2pi/3) will have the equivalent rectangular coordinates x= -7 * cos (2pi/3) = -7 * (-1/2) = 7/2 y= -7 * sin(2pi/3) = -7 * sqrt(3) / 2

OpenStudy (anonymous):

oh thank you very much :)